# Quantifying Lexical Semantic Shift via Unbalanced Optimal Transport

### v1

@@ Acknowledgments
This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, 23K24910, JST CREST JPMJCR21N3, JST SPRING JPMJSP2110.
@@ Appendix A Further Illustrative Examples
Figure 8: Re-rendering of the t-SNE visualization for the word record in Fig. 1. The color of each instance represents the gold sense in the DWUG dataset.
@@
Observing the visualizations of the five target words shown in Fig. 7, it can be seen that both SUS and LDR roughly reflect the sense clusters appropriately. However, a closer inspection reveals differences between the two methods. For the target word record, referring to the gold senses provided by the DWUG dataset in Fig. 8, it is evident that SUS reflects the clusters more accurately than LDR. In particular, in the rightmost part of the cluster corresponding to the sense [information], there is a region where circles and boxes are mixed. Compared to SUS, LDR more clearly recognizes this as distinct clusters. For the target word bar, the sense [music] newly emerged in the modern corpus. SUS correctly identifies this increase in usage frequency, whereas LDR incorrectly identifies it as a frequency decrease. However, such detailed observations have not been sufficiently discussed, and caution is needed when drawing conclusions from these differences.
@@ Appendix B Details of Log-Density Ratio
In our experiments, we directly computed LDR from Eq. (8) of Nagata et al. (2023). Specifically, we calculated the approximate maximum likelihood estimator of the concentration parameter $\kappa$ for the von Mises-Fisher (vMF) distribution for the old and modern corpora using Eq. (5) therein, and then derived the two corresponding density functions of the vMF distribution using vonmises_fisher in scipy.
@@ Appendix E Details of XL-LEXEME
In this appendix, we describe how XL-LEXEME [^7] https://huggingface.co/pierluigic/xl-lexeme (Cassotti et al., 2023) calculates the embeddings of target words. XL-LEXEME is a model obtained by fine-tuning XLM-RoBERTa (Conneau et al., 2020) on Word-in-Context (Pilehvar and Camacho-Collados, 2019) and it has 561M parameters. In the DWUG dataset, each usage instance of a target word is accompanied by positional information that specifies the position of the target word within its context. XL-LEXEME takes this context and positional information as input to compute the embedding for the target word. Specifically, special tokens are inserted before and after the target word in the context, and the embedding of the target word is computed as the mean of the embeddings of all tokens in the context.
@@ Results.
Table 2 shows the Spearman rank correlation between the gold graded change scores $\tau^{*}$ and the change score produced by each method. The estimator based on SUS outperforms the sense-based method, providing more accurate predictions of changes in word usage frequency for the sense of each instance.

### v2

@@ Acknowledgments
We would like to thank the anonymous reviewers for their helpful comments and suggestions. This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, 23K24910, JST FOREST JPMJFR2331, and JST CREST JPMJCR21N3.
@@ Appendix A Further Illustrative Examples
As described in Section 4.4, Fig. 7 indicates that both SUS and LDR generally capture the sense clusters reasonably well. However, a more detailed examination highlights distinctions between the two approaches. For the target word record, when compared against the gold senses from the DWUG dataset shown in Fig. 5, SUS appears to align more closely with the actual clusters than LDR. Specifically, in the rightmost portion of the cluster representing the sense [information], there is a region where circles and squares overlap. LDR distinguishes this area as separate clusters more clearly than SUS does. For the target word bar, the sense [music] newly emerged in the modern corpus. SUS correctly identifies this increase in usage frequency, whereas LDR incorrectly identifies it as a frequency decrease. However, such detailed observations have not been sufficiently discussed, and caution is needed when drawing conclusions from these differences.
@@ Appendix B Details of Log-Density Ratio
In our experiments, we directly computed LDR from Eq. (8) of Nagata et al. (2023). Specifically, we calculated the approximate maximum likelihood estimator (MLE) of the concentration parameter $\kappa$ for the von Mises-Fisher (vMF) distribution for the old and modern corpora as follows: Let $\{\bm{u}_{i}\}_{i=1}^{m}\subset\mathbb{R}^{d}$ denote the set of embeddings in the old corpus and define $\tilde{\bm{u}}_{i}=\bm{u}_{i}/\|\bm{u}_{i}\|$. Assume that $\{\tilde{\bm{u}}_{i}\}_{i=1}^{m}$ follow a vMF distribution with mean direction parameter $\bm{\mu}_{S}$ and concentration parameter $\kappa_{S}$. Let $\ell=\left\|\frac{1}{m}\sum_{i}\tilde{\bm{u}}_{i}\right\|$. Then, the maximum likelihood estimator (MLE) of $\bm{\mu}_{S}$ is given by $\frac{1}{\ell m}\sum_{i}\tilde{\bm{u}}_{i}$. While the MLE of $\kappa_{S}$ does not have a closed-form solution, Banerjee et al. (2005) provide the following approximation: $\kappa_{S}\approx\frac{\ell(d-\ell^{2})}{1-\ell^{2}}$. After calculating the MLEs, we derived the corresponding vMF density function using vonmises_fisher in scipy. We applied the same procedure to the modern embeddings $\{\bm{v}_{j}\}_{j=1}^{n}$ by computing their normalized forms $\tilde{\bm{v}}_{j}=\bm{v}_{j}/\|\bm{v}_{j}\|$ and estimating the parameters $\bm{\mu}_{T}$ and $\kappa_{T}$ of the associated vMF distribution.
@@ Appendix E Details of XL-LEXEME
We describe how XL-LEXEME [^9] https://huggingface.co/pierluigic/xl-lexeme (Cassotti et al., 2023) calculates the embeddings of target words. XL-LEXEME is a model obtained by fine-tuning XLM-RoBERTa (Conneau et al., 2020) on Word-in-Context (Pilehvar and Camacho-Collados, 2019) and it has 561M parameters. In the DWUG dataset, each usage instance of a target word includes positional information specifying the position of the target word within its context. XL-LEXEME takes this context and positional information as input to compute the embedding for the target word. Specifically, special tokens are inserted before and after the target word in the context, and the embedding of the target word is computed as the mean of the embeddings of all tokens in the context.
@@ Appendix F Details of Experiment in Section 6
Method Approach Stable (25) Changed (21) Overall (46)
$\tau_{\mathrm{SUS}}$ SUS-based 0.21 0.61 0.46
$\tau_{\mathrm{LDR}}$ form-based 0.10 0.34 0.40
$\tau_{\mathrm{WiDiD}}$ sense-based 0.15 0.41 0.31
Table 9: Performance of methods for predicting the instance-level change score. Overall corresponds to ‘Instance-level’ in Table 2.
Method Approach Stable (25) Changed (21) Overall (46)
$\tau_{\mathrm{SUS}}$ SUS-based 0.70 0.85 0.83
$\tau_{\mathrm{LDR}}$ form-based 0.64 0.38 0.70
$\tau_{\mathrm{WiDiD}}$ sense-based 0.54 0.93 0.84
Table 10: Performance of methods for predicting the sense-level change score. Overall corresponds to ‘Sense-level’ in Table 2.
@@ Results.
In the DWUG dataset, each target word is annotated with a label indicating whether its meaning has changed, based on old and modern SFDs. We divide the test set words according to this change annotation and report the instance-level and sense-level performances for each group in Table 9 and Table 10, respectively. The overall score is computed without splitting the test set and corresponds to Table 2. Hyperparameter tuning was performed using the validation set without splitting words into stable or changed categories. Performance evaluation was then conducted separately for stable and changed words in the test set.
For stable words, instance-level evaluation performs quite poorly across all methods. In stable words, semantic change is minimal, but the values of each metric still fluctuate, and even small variations in these values can result in significant changes in rankings. Therefore, instance-level rank correlation is not the most appropriate evaluation metric in this context. In any way, SUS demonstrates superior instance- and sense-level performance compared to LDR not only overall but also within both the stable and changed word categories.
@@
Appendix H Experiments on Another Dataset
Method Approach Spearman
$f_{\mathrm{SUS}}$ SUS-based 0.61
$f_{1}$ SUS-based 0.60
$f_{2}$ SUS-based 0.60
$f_{3}$ SUS-based 0.61
$f_{\mathrm{OT}}$ form-based 0.61
$f_{\mathrm{APD}}$ form-based 0.61
$f_{\mathrm{LDR}}$ form-based 0.36
$f_{\mathrm{WiDiD}}$ sense-based 0.41
$f_{\mathrm{APDP}}$ sense-based 0.52
Table 13: Performance of methods for measuring the magnitude of word-level semantic change in the Spanish DWUG dataset.
Method Approach Spearman
$g_{\mathrm{SUS}}$ SUS-based 0.36
$g_{1}$ SUS-based 0.45
$g_{\mathrm{vMF}}$ form-based 0.41
$g_{\mathrm{LDR}}$ form-based 0.31
$g_{\mathrm{WiDiD}}$ sense-based 0.33
Table 14: Performance of methods for measuring word-level changes in semantic scope in the Spanish DWUG dataset.
Since the SUS-based metric evaluated in Section 7 may be overfitting to the English DWUG dataset (DWUG EN), we conduct an evaluation experiment for word-level semantic change detection using the Spanish DWUG dataset Zamora-Reina et al. (2022) (DWUG ES) as an additional dataset to verify the effectiveness of SUS. The experimental setup including hyperparameter tuning follow the same procedure as in Section 7. The search scope for the hyperparameter is the same as in the experiment with DWUG EN.
Quantifying the magnitude of word-level semantic change.
Table 13 shows the Spearman rank correlation between the gold change score $f^{*}$ and the change score produced by each method. The SUS-based approach demonstrates performance comparable to the SOTA method $f_{\mathrm{APD}}$, suggesting that SUS is also effective in measuring the degree of semantic change in DWUG ES.
Quantifying the word-level change in semantic scope.
Table 14 shows the Spearman rank correlation between the gold change score $g^{*}$ and the change score calculated by each method. For $g_{\mathrm{SUS}}$, when compared to the baselines, the ranking of results is consistent with that in DWUG EN, and for $g_{1}$, it outperforms $g_{\mathrm{vMF}}$. This suggests that SUS is also effective in capturing changes in semantic scope in DWUG ES.